# Current address: ]College of William and Mary, Williamsburg, Virginia 23187-8795

### v1

@@ I Introduction
The study of nucleon structure has been an active field of research ever since the discovery that nucleons were composite [1, 2, 3]. It is now well known that the nucleon, like all hadrons, is made of partons (quarks, antiquarks, and gluons) and that its structure is dominantly ruled by the strong nuclear interaction described by quantum chromodynamics (QCD) [4, 5]. However, the behavior of the QCD coupling constant $\alpha_{s}$ [6] determines that our understanding of nucleon structure at long distance is not as developed as that at short distance. At high energy-momentum (short distance) typically well above the GeV scale, $\alpha_{s}\ll 1$, which allows us to treat QCD perturbatively. Presently, the approximations of observables are typically known to order $\alpha_{s}^{3}$, and some to $\alpha_{s}^{5}$. On the other hand, at low energy-momentum, $\alpha_{s}\gtrsim 1$, and a perturbative expansion in $\alpha_{s}$ becomes inapplicable. The problem is further complicated by the rise of complex phenomena, e.g. the confinements of partons, the emergence of hadronic degrees of freedom or the spontaneous breaking of QCD’s near $SU(3)_{L}\times SU(3)_{R}$ chiral symmetry. In this domain, due to the lack of a mature analytic technique based on the QCD Lagrangian, effective theories or models are often used. In particular, chiral effective field theory ($\chi$EFT) [7], based on the observed spontaneous breaking of the chiral symmetry in hadronic states, has successfully predicted hadronic observables at low energies.
@@
The principle of $\chi$EFT is that it describes the nucleon in terms of effective hadronic degrees of freedom instead of the fundamental quark and gluon fields. It is completely non-perturbative in terms of $\alpha_{s}$ because it employs series whose expansion parameters are based e.g. on the pion mass, the nucleon-$\Delta_{1232}$ mass gap or the characteristic energy scale of the chiral symmetry breaking. However, while leading-order (LO) calculations of $\chi$EFT are relatively tractable, next-to-leading-order (NLO) and higher order calculations are much more involved. An important experimental task is therefore to provide precise hadron structure data that can test, and potentially guide, $\chi$EFT calculations and other approaches to non-perturbative QCD. Such data must cover the domain of low momentum transfer squared ($Q^{2}$) where $\chi$EFT applies, typically below a few tenths of GeV2. Especially important in this context are low $Q^{2}$ nucleon spin structure data [8, 9, 10] which are challenging to obtain experimentally but are crucial for the development of $\chi$EFT [11]. One such experimental endeavor is that of the Jefferson Lab (JLab) Experimental Group EG4, constituted of experiments E03006 and E05111 that used the CEBAF Large Acceptance Spectrometer (CLAS) in Hall B. EG4 was designed to test $\chi$EFT predictions by precisely measuring the longitudinal spin structure function $g_{1}$ and its moments for the proton and the deuteuron down to $Q^{2}\simeq 0.01$ GeV2. These low $Q^{2}$ data, reported in two letters [12, 13], provided benchmarks for $\chi$EFT calculations. This document is the archival article that presents in detail the EG4 experiments, data analysis procedure for the inclusive scattering channel, results on the proton and the deuteron, and followup analysis and results on the neutron spin structure functions. Additional results on the target- and double-spin asymmetries of pion electroproduction are available [14], but these were focused on the semi-inclusive channel and are not included here.
@@
The article is arranged as follows: In Section II we present the formalism of unpolarized and polarized electron scattering. We introduce the nucleon structure functions, their moments, and polarized sum rules, which are relevant not only for testing $\chi$EFT but also to form observables that have no known method of direct measurement. In Section III we describe EG4 and the special features that distinguish it from other CLAS experiments. In Section IV we describe the analysis procedure of extracting the polarized yield differences from the data. In Section V we describe the Monte-Carlo simulation, the normalization of the data using the elastic scattering yield, and the extraction of the structure function $g_{1}$. In Section VI we present results on $g_{1}$, $A_{1}F_{1}$ of the proton and the deuteron, followed by the procedure to extract the neutron $g_{1}$ results. We present results for and discussions of all the moments in Section VII. We then summarize, conclude, and provide perspectives in Section VIII.
@@ II.4 Sum Rules and Moments of Structure Functions
In this section we present the Bjorken, GDH, generalized GDH, and generalized polarizability sum rules, which all involve moments of polarized structure functions or, equivalently, $\sigma_{TT}$ and $\sigma_{LT}$. For reviews of nucleon spin structure and additional sum rules, see Refs. [9, 10].
@@ II.4.1 Bjorken Sum Rule
The Bjorken sum rule, derived for $Q^{2}\to\infty$ using current algebra and isospin symmetry [18, 19], predates QCD and is not a QCD prediction. However, its generalization to include the $Q^{2}$-dependence appearing at finite $Q^{2}$ [20, 21, 22] stems from pQCD. Studying the sum rule in the pQCD domain ($Q^{2}\gtrsim 1$ GeV2) therefore tests whether QCD correctly describes the strong force when spin degrees-of-freedom are explicit. The Bjorken sum has been actively measured in this domain at CERN [23, 24, 25, 26, 27, 28], DESY [29, 30, 31], JLab [32], and SLAC [33, 34, 35, 36, 37, 38, 39, 40, 41, 42], with measurement agreeing to better than 10% with the sum rule expectation [10].
@@ II.4.2 Gerasimov-Drell-Hearn (GDH) Sum Rule
where the bars and $1^{-}$ indicate that the $x=1$ elastic contribution (quasi-elastic for nuclear targets) is excluded. In the quark-parton model, this integral gives the quark polarization within the nucleon weighted by their electric charge squared, and provides one of the ingredients for Ellis-Jaffe sum rule [72] and Bjorken sum rule. For the generalized form of the GDH sum rule, Eq. (33), the amplitude $\bar{S}_{1}(0,Q^{2})$ can be calculated in $\chi$EFT if $Q^{2}$ is sufficiently low [73, 74, 75, 76, 77, 53, 78, 79, 54].
@@ III.1 The Polarized Electron Beam and Beam Monitoring
The EG4 experiment used a longitudinally polarized beam. The beam helicity was flipped at 30 Hz, taking the quartet structure of either + - - + or - + + -, with the first helicity window selected from a pseudo-random sequence. The beam polarization was regularly measured by a Möller polarimeter set at the entrance of the Hall. It consisted of detectors, transport and polarizing magnets, and a 25 $\mu$m thick Permendur foil (49% Fe, 49% Co, 2% Va) acting as a polarized electron target and oriented at ±$\pm 20^{\circ}$ with respect to the beamline. A pair of 120-G Helmholtz coils thermally polarized the foil longitudinally to 7.5%. For each beam polarization measurement, the foil was inserted in the beam for about 30 minutes. The scattered (recoil) electrons were transported to the detectors (two lead-glass blocks) by two quadrupoles. The Möller asymmetry was measured by detecting the two electrons in coincidence and compared to the well-known theoretical expectation to obtain the beam polarization. A 30-min measurement provided a $\approx$1statistical precision. The beam polarization for EG4 was typically around 86% [14], determined with a $\approx$2.5systematic uncertainty. Two of the Möller measurements were cross-checked with a Mott polarimeter set in the CEBAF injector that measured the polarization of 5 MeV electrons [99]. These direct measurements were used to check the stability of the beam polarization, while for the data analysis the product of beam and target polarization $P_{b}P_{t}$ was directly determined in $ep$ elastic or $e$D quasi-elastic scattering, see Sections V-A and V-D.
@@ IV Data Analysis
The analysis presented in this section lead to the extraction of both the proton and the deuteron spin structure functions $g_{1}$, $A_{1}F_{1}$, and their moments, previously reported in Refs. [12, 13]. We will present the general method (Section IV-A), event selection criteria and kinematic corrections (Section IV-B), particle identification performance and efficiency analysis of the new CC (Section IV-C), and background (Section IV-D).
@@ IV.1 Overview of the Analysis Methods
After event selection and kinematic correction (see next two sections), the data were divided into the $Q^{2}$ bins shown in Table I, with bin limits determined logarithmically (the upper edge of each bin $Q^{2}_{\mathrm{max}}$ is $10^{1/13}$ of the lower edge $Q^{2}_{\mathrm{min}}$). In practice, proton data exist from $Q^{2}$ bin 1 to 25, and deuteron data from bin 4 to 25.
@@ IV.3 Particle Identification and Cherenkov Efficiency
Once we extracted the CC efficiency and its uncertainty map for given $I_{torus}/p$ range and $(\theta_{vtx},\phi_{DC1})$ bin, we matched it to the $n_{\rm p.e.}$ map for the same bin. Then we fit the values of the efficiency as a function of $n_{\rm p.e.}$ using the Poisson function form, see Fig. 8. We studied the efficiency fit for both individual $I_{torus}/p$ bins as well as all bins combined. (When combining bins, we combined $N_{1}$ and $N_{2}$ first and then evaluated the efficiency and its uncertainty.) We found the fit to achieve good results for all cases, and the fit results for individual $I_{torus}/p$ bins were consistent with those achieved by combining all bins, see Fig. 8. These results were used in the simulation to account for the CC efficiency, see Section V.
Figure 8: CC efficiency vs. the number of photoelectrons $n_{\rm p.e.}$ for the $I_{torus}/2250$A$/p$ (GeV)$=(0.6,0.7)$ bin (top) and for all $I_{torus}/p$ bins combined (bottom) for 2.3 GeV data taken on the NH3 top long cell. The efficiencies and error bars were evaluated using Eqs. (40-42). The fit results for the $I_{torus}/2250$A$/p$(GeV)$=(0.6,0.7)$ bin is $\eta_{cc}=0.986-e^{-0.062(x-0.44)}$ (dashed curve), and for the combined results is $\eta_{cc}=0.987-e^{-0.040(x-18.8)}$ (solid curve). When making these plots, all good electron cuts were applied, thus there is no entry below $n_{\rm p.e.}=2$ by definition. In practice, there were very few or no event below $n_{\rm p.e.}=5$.
@@ IV.4 Background Contamination
For inelastic scattering processes, charged pions can be misidentified as electrons. Similarly, neutral pion decay can produce photons and subsequently electrons via the pair production process. Contamination from these two background processes were studied using charged pion yields, and positron yields taken with the in-bending torus configuration. The pion contamination was found to be less than 1%, and pair production yield was on the order of $10^{-3}$ for all kinematics. We did not correct these backgrounds, but instead incorporated them into the systematic uncertainties. An additional background is the electron scattering off the 15N in the target. The 15N is slightly polarized, and thus we discuss this background along with its correction in Section V-H as part of the systematic uncertainties.
@@ V.1 Simulation of (Quasi-)Elastic Polarized Yield Differences
A simulation for $ep$ elastic and $e$D quasi-elastic scattering events was performed using GSIM, the standard CLAS simulation program, combined with an event generator for (quasi-)elastic scattering. Events were generated for ranges of $\theta=(5^{\circ},45^{\circ})$ and $\phi=(250^{\circ},325^{\circ})$. The $\theta$ and $\phi$ ranges ensured the whole Sector 6 was covered, while keeping a reasonable generator efficiency.
@@
The simulated data were then passed through the GSIM Post Processor (GPP) to account for smearing factors in the detectors, and RECSIS, the standard CLAS simulation program for event reconstruction. Detailed inputs to the simulation, GPP, and corrections applied in the reconstruction are given below.
@@ V.3 Smearing Factor for DC Resolution
where $x$ is the DC wire number and $p_{0,1,2}$ are fit parameters, see Table II. The smearing factor starts from a value greater than 1 for low wire numbers (small scattering angle) to approximately 1.0 for high wire numbers (large scattering angle). The large smearing factor for small scattering angle was possibly due to limits in the knowledge of the magnetic field of CLAS in that region.
@@ V.4 Comparison of Elastic Simulation with Data
Figure 9 shows a representative bin for the comparison between data and simulation for elastic $ep$ scattering. Data from low $Q^{2}$ bins were not used for normalization because of the 15N elastic contamination (see Section V-H). Both the energy loss correction and the DC smearing were applied to the simulation.
@@
Careful studies were done for the elastic simulation quality for individual $Q^{2}$ bins. Gaussian fits were used to compare the peak location and width of simulation with data to ensure that the simulated elastic peak agrees almost perfectly with the data (and the peak position agrees with the proton mass). Bins where the simulated peak position or shape deviate from data were excluded from the normalization input. This happened only at very low $Q^{2}$ bins due to 15N contamination (see Section V-H), and at very high $Q^{2}$ bins due to sensitivity to the low-efficiency area of the DC as shown in Fig. 7. After rejecting the very low and very high $Q^{2}$ bins, we normalized the simulated elastic peak to that from data, summed over all $Q^{2}$ bins. The normalization for the deuteron data proceeded along the same lines, using the simulation of the quasi-elastic peak instead.
The uncertainty in the elastic peak normalization included the uncertainty in the proton form factor input [111], which was found to be less than $2\%$ at 1.1 GeV and increased slightly for higher beam energies. The uncertainty due to radiative corrections was found to be $\approx 1\%$ for all energies. In addition, we compared [104] with a recent proton form factor fit [112] that accounted for the uncertainty in the proton radius, and found the effect to be less than 1%. Overall, we estimated the uncertainty in elastic normalization to be $(1+E_{b})\%$ with $E_{b}$ the beam energy in GeV. For the deuteron normalization based on the quasi-elastic peak, we estimated a total uncertainty of 10% for each beam energy, accounting for form factor uncertainties, the observed variation of the measured yield to simulated yield ratio with $Q^{2}$, and uncertainties in the folding and radiative correction procedure. These uncertainties were taken into account in our estimate of the systematic uncertainties (see Section V-I).

### v2

@@ I Introduction
The study of nucleon structure has been an active field of research ever since the discovery that nucleons were composite [1, 2, 3]. It is now well known that the nucleon, like all hadrons, is made of partons (quarks, antiquarks, and gluons) and that its structure is dominantly ruled by the strong nuclear interaction described by quantum chromodynamics (QCD) [4, 5]. However, the behavior of the QCD coupling constant $\alpha_{s}$ [6, 7, 8] determines that our understanding of nucleon structure at long distance is not as developed as that at short distance. At high energy-momentum (short distance) typically well above the GeV scale, $\alpha_{s}\ll 1$, which allows us to treat QCD perturbatively. Presently, the approximations of observables are typically known to order $\alpha_{s}^{3}$, and some to $\alpha_{s}^{5}$. On the other hand, at low energy-momentum, $\alpha_{s}\gtrsim 1$, and a perturbative expansion in $\alpha_{s}$ becomes inapplicable. The problem is further complicated by the rise of complex phenomena, e.g. the confinements of partons, the emergence of hadronic degrees of freedom or the spontaneous breaking of QCD’s near $SU(3)_{L}\times SU(3)_{R}$ chiral symmetry. In this domain, due to the lack of a mature analytic technique based on the QCD Lagrangian, effective theories or models are often used. In particular, chiral effective field theory ($\chi$EFT) [9], based on the observed spontaneous breaking of the chiral symmetry in hadronic states, has successfully predicted hadronic observables at low energies.
@@
The principle of $\chi$EFT is that it describes the nucleon in terms of effective hadronic degrees of freedom instead of the fundamental quark and gluon fields. It is completely non-perturbative in terms of $\alpha_{s}$ because it employs series whose expansion parameters are based e.g. on the pion mass, the nucleon-$\Delta_{1232}$ mass gap or the characteristic energy scale of the chiral symmetry breaking. However, while leading-order (LO) calculations of $\chi$EFT are relatively tractable, next-to-leading-order (NLO) and higher order calculations are much more involved. An important experimental task is therefore to provide precise hadron structure data that can test, and potentially guide, $\chi$EFT calculations and other approaches to non-perturbative QCD. Such data must cover the domain of low momentum transfer squared ($Q^{2}$) where $\chi$EFT applies, typically below a few tenths of GeV2. Especially important in this context are low $Q^{2}$ nucleon spin structure data [10, 11, 11, 12, 13, 14, 15, 16, 17, 18] which are challenging to obtain experimentally but are crucial for the development of $\chi$EFT [19]. One such experimental endeavor is that of the Jefferson Lab (JLab) Experimental Group EG4, constituted of experiments E03006 and E05111 that used the CEBAF Large Acceptance Spectrometer (CLAS) in Hall B [20]. EG4 was designed to test $\chi$EFT predictions by precisely measuring the longitudinal spin structure function $g_{1}$ and its moments for the proton and the deuteuron down to $Q^{2}\simeq 0.01$ GeV2. These low $Q^{2}$ data, reported in two letters [21, 22], provided benchmarks for $\chi$EFT calculations. This document is the archival article that presents in detail the EG4 experiments, data analysis procedure for the inclusive scattering channel, results on the proton and the deuteron, and followup analysis and results on the neutron spin structure functions. Additional results on the target- and double-spin asymmetries of pion electroproduction are available [23], but these were focused on the semi-inclusive channel and are not included here.
@@
The article is arranged as follows: In Section II we present the formalism of unpolarized and polarized electron scattering. We introduce the nucleon structure functions, their moments, and polarized sum rules, which are relevant not only for testing $\chi$EFT but also to form observables that have no known method of direct measurement. In Section III we describe EG4 and the special features that distinguish it from other CLAS experiments. In Section IV we describe the analysis procedure of extracting the polarized yield differences from the data. In Section V we describe the Monte-Carlo simulation, the normalization of the data using the elastic scattering yield, and the extraction of the structure function $g_{1}$. In Section VI.1 we present results on $g_{1}$, $A_{1}F_{1}$ of the proton and the deuteron, followed by the procedure to extract the neutron $g_{1}$ results. We present results for and discussions of all the moments in Section-VII. We then summarize, conclude, and provide perspectives in Section VIII.
@@ II.4 Sum Rules and Moments of Structure Functions
In this section we present the Bjorken, GDH, generalized GDH, and generalized polarizability sum rules, which all involve moments of polarized structure functions or, equivalently, $\sigma_{TT}$ and $\sigma_{LT}$. For reviews of nucleon spin structure and additional sum rules, see Refs. [12, 14, 15, 17].
@@ II.4.1 Bjorken Sum Rule
The Bjorken sum rule, derived for $Q^{2}\to\infty$ using current algebra and isospin symmetry [27, 28], predates QCD and is not a QCD prediction. However, its generalization to include the $Q^{2}$-dependence appearing at finite $Q^{2}$ [29, 30, 31] stems from pQCD. Studying the sum rule in the pQCD domain ($Q^{2}\gtrsim 1$ GeV2) therefore tests whether QCD correctly describes the strong force when spin degrees-of-freedom are explicit. The Bjorken sum has been actively measured at CERN [32, 33, 34, 35, 36, 37, 38, 39, 40], DESY [41], JLab [42, 43, 44], and SLAC [45, 46, 47, 48, 49, 50], with measurement agreeing to better than 10% with the sum rule expectation [17]. Measurements performed with the higher-energy accelerators, CERN and SLAC, provide data at larger $Q^{2}$ values, offering extensive coverage of the low-$x$ part of the Bjorken sum. Measurements at the lower-energy facilities, DESY and JLab, covered the smaller $Q^{2}$ range while overlapping with the SLAC data. Due to their lower energy, DESY and JLab have a limited low-$x$ reach, which is supplemented by extrapolating the low-$x$ data obtained from CERN and SLAC.
@@ II.4.2 Gerasimov-Drell-Hearn (GDH) Sum Rule
where the bars and $1^{-}$ indicate that the $x=1$ elastic contribution (quasi-elastic for nuclear targets) is excluded. In the quark-parton model, this integral gives the quark polarization within the nucleon weighted by their electric charge squared, and provides one of the ingredients for Ellis-Jaffe sum rule [78] and Bjorken sum rule. For the generalized form of the GDH sum rule, Eq. (33), the amplitude $\bar{S}_{1}(0,Q^{2})$ can be calculated in $\chi$EFT if $Q^{2}$ is sufficiently low [79, 80, 81, 82, 83, 59, 84, 85, 60]. The generalized GDH integrals have been measured over an extensive $Q^{2}$ range and on proton, neutron, deuteron and 3He by CERN [86, 32, 33, 34, 35, 36, 37, 87, 88, 38, 39, 40, 89], DESY [90, 91, 41], JLab [92, 93, 94, 95, 96, 51, 97, 98, 99, 100, 101, 102, 103, 21, 22, 71, 104], and SLAC [105, 106, 107, 45, 46, 47, 108, 48, 49, 50]. CERN and SLAC provide the higher $Q^{2}$ coverage, with extensive contributions to the moments from the low-$x$ region. JLab data cover the lowest $Q^{2}$ domain. All accelerators provide data points in the intermediate $Q^{2}$ region of a few GeV or less, including the COMPASS experiment [87, 88, 38, 39, 40, 89] for CERN, E143 [105, 106, 107, 45, 46] for SLAC and HERMES for DESY [90, 91, 41].
@@ III.1 The Polarized Electron Beam and Beam Monitoring
The EG4 experiment used a longitudinally polarized beam. The beam helicity was flipped at 30 Hz, taking the quartet structure of either + - - + or - + + -, with the first helicity window selected from a pseudo-random sequence. The beam polarization was regularly measured by a Möller polarimeter set at the entrance of the Hall. It consisted of detectors, transport and polarizing magnets, and a 25 $\mu$m thick Permendur foil (49% Fe, 49% Co, 2% Va) acting as a polarized electron target and oriented at ±$\pm 20^{\circ}$ with respect to the beamline. A pair of 120-G Helmholtz coils thermally polarized the foil longitudinally to 7.5%. For each beam polarization measurement, the foil was inserted in the beam for about 30 minutes. The scattered (recoil) electrons were transported to the detectors (two lead-glass blocks) by two quadrupoles. The Möller asymmetry was measured by detecting the two electrons in coincidence and compared to the well-known theoretical expectation to obtain the beam polarization. A 30-min measurement provided a $\approx$1statistical precision. The beam polarization for EG4 was typically around 86% [23], determined with a $\approx$2.5systematic uncertainty. Two of the Möller measurements were cross-checked with a Mott polarimeter set in the CEBAF injector that measured the polarization of 5 MeV electrons [123]. These direct measurements were used to check the stability of the beam polarization, while for the data analysis the product of beam and target polarization $P_{b}P_{t}$ was directly determined in $ep$ elastic or $e$D quasi-elastic scattering, see Sections V.1 and V.4.
@@ IV Data Analysis
The analysis presented in this section lead to the extraction of both the proton and the deuteron spin structure functions $g_{1}$, $A_{1}F_{1}$, and their moments, previously reported in Refs. [21, 22]. We will present the general method (Section IV.1), event selection criteria and kinematic corrections (Section IV.2), particle identification performance and efficiency analysis of the new CC (Section IV.3), and background (Section IV.4).
@@ IV.1 Overview of the Analysis Methods
After event selection and kinematic correction (see next two sections), the data were divided into the $Q^{2}$ bins shown in Table 1, with bin limits determined logarithmically (the upper edge of each bin $Q^{2}_{\mathrm{max}}$ is $10^{1/13}$ of the lower edge $Q^{2}_{\mathrm{min}}$). In practice, proton data exist for all 25 $Q^{2}$ bins, and deuteron data from bin 3 to 24. Our neutron results (as presented in Section VI.2) were extracted from combining proton with deuteron data, and thus also exist from bins 3 to 24 only.
@@ IV.3 Particle Identification and Cherenkov Efficiency
Once we extracted the CC efficiency and its uncertainty map for given $I_{torus}/p$ range and $(\theta_{vtx},\phi_{DC1})$ bin, we matched it to the $n_{\rm p.e.}$ map for the same bin. Then we fit the values of the efficiency as a function of $n_{\rm p.e.}$ using the Poisson function form, see Fig. 8 top panel. We studied the efficiency fit for both individual $I_{torus}/p$ bins as well as all bins combined. (When combining bins, we combined $N_{1}$ and $N_{2}$ first and then evaluated the efficiency and its uncertainty.) We found the fit to achieve good results for all cases, and the fit results for individual $I_{torus}/p$ bins were consistent with those achieved by combining all bins, see Fig. 8 bottom panel. These results were used in the simulation to account for the CC efficiency, see Section V.
Figure 8: CC efficiency (open circles) vs. the number of photoelectrons $n_{\rm p.e.}$ for the $I_{torus}/2250$A$/p$ (GeV)$=(0.6,0.7)$ bin (top) and for all $I_{torus}/p$ bins combined (bottom) for 2.3 GeV data taken on the NH3 top long cell. The efficiencies and error bars were evaluated using Eqs. (40-42). The fit results for the $I_{torus}/2250$A$/p$(GeV)$=(0.6,0.7)$ bin is $\eta_{cc}=0.986-e^{-0.062(x-0.44)}$ (dashed curve), and for the combined results is $\eta_{cc}=0.987-e^{-0.040(x-18.8)}$ (solid curve). When making these plots, all good electron cuts were applied, thus there is no entry below $n_{\rm p.e.}=2$ by definition. Note that even after combining all bins, the region with $n_{\rm p.e.}\leqslant 6$ has low statistics and relatively large uncertainties.
@@ IV.4 Background Contamination
For inelastic scattering processes, charged pions can be misidentified as electrons. Similarly, neutral pion decay can produce photons and subsequently electrons via the pair production process. Contamination from these two background processes were studied using charged pion yields, and positron yields taken with the in-bending torus configuration. The pion contamination was found to be less than 1%, and pair production yield was on the order of $10^{-3}$ for all kinematics. We did not correct these backgrounds, but instead incorporated them into the systematic uncertainties. An additional background is the electron scattering off the 15N in the target. The 15N is slightly polarized, and thus we discuss this background along with its correction in Section V.8 as part of the systematic uncertainties.
@@ V.1 Simulation of (Quasi-)Elastic Polarized Yield Differences
A simulation for $ep$ elastic and $e$D quasi-elastic scattering events was performed using GSIM [128, 129], the standard CLAS simulation program, combined with an event generator for (quasi-)elastic scattering. Events were generated for ranges of $\theta=(5^{\circ},45^{\circ})$ and $\phi=(250^{\circ},325^{\circ})$. The $\theta$ and $\phi$ ranges ensured the whole Sector 6 was covered, while keeping a reasonable generator efficiency.
@@
The simulated data were then passed through the GSIM Post Processor (GPP) [128] to account for smearing factors in the detectors, and RECSIS [137], the standard CLAS simulation program for event reconstruction. Detailed inputs to the simulation, GPP, and corrections applied in the reconstruction are given below.
@@ V.3 Smearing Factor for DC Resolution
where $x$ is the DC wire number and $p_{0,1,2}$ are fit parameters, see Table 2. The smearing factor starts from a value greater than 1 for low wire numbers (small scattering angle) to approximately 1.0 for high wire numbers (large scattering angle). The large smearing factor for small scattering angle was possibly due to limits in the knowledge of the magnetic field of CLAS in that region.
@@ V.4 Comparison of Elastic Simulation with Data
Figure 9 shows a representative bin for the comparison between data and simulation for elastic $ep$ scattering. Data from low $Q^{2}$ bins were not used for normalization because of the 15N elastic contamination (see Section V.8). Both the energy loss correction and the DC smearing were applied to the simulation.
@@
Careful studies were done for the elastic simulation quality for individual $Q^{2}$ bins. Gaussian fits were used to compare the peak location and width of simulation with data to ensure that the simulated elastic peak agrees almost perfectly with the data (and the peak position agrees with the proton mass). Bins where the simulated peak position or shape deviate from data were excluded from the normalization input. This happened only at very low $Q^{2}$ bins due to 15N contamination (see Section V.8), and at very high $Q^{2}$ bins due to sensitivity to the low-efficiency area of the DC as shown in Fig. 7. After rejecting the very low and very high $Q^{2}$ bins, we normalized the simulated elastic peak to that from data, summed over all $Q^{2}$ bins. The normalization for the deuteron data proceeded along the same lines, using the simulation of the quasi-elastic peak instead.
The uncertainty in the elastic peak normalization included the uncertainty in the proton form factor input [138], which was found to be less than $2\%$ at 1.1 GeV and increased slightly for higher beam energies. The uncertainty due to radiative corrections was found to be $\approx 1\%$ for all energies. In addition, we compared [130] with a recent proton form factor fit [139] that accounted for the uncertainty in the proton radius, and found the effect to be less than 1%. Overall, we estimated the uncertainty in elastic normalization to be $(1+E_{b})\%$ with $E_{b}$ the beam energy in GeV. For the deuteron normalization based on the quasi-elastic peak, we estimated a total uncertainty of 10% for each beam energy, accounting for form factor uncertainties, the observed variation of the measured yield to simulated yield ratio with $Q^{2}$, and uncertainties in the folding and radiative correction procedure. These uncertainties were taken into account in our estimate of the systematic uncertainties (see Section V.9).